# ISAC channel simulator GBSM

### v1

@@ Abstract
Integrated Sensing and Communication (ISAC) is one of the key technologies in 6G, and related research and standardization efforts are progressing vigorously. Wireless channel simulation is the cornerstone for the evaluation and optimization of wireless communication technologies. This paper proposes a design and implementation method for an ISAC channel simulation based on a Geometry-Based Stochastic Model (GBSM) simulation framework. First, we introduce the progress of 3GPP ISAC channel standardization and the key topics of discussion. Second, addressing the current lack of a standardized ISAC channel simulation framework, we propose a cascaded ISAC channel simulation framework based on GBSM, leveraging our team’s related measurements, analyses, and proposal results. Based on this framework, we develop and design the ISAC channel simulator BUPTCMCC-6G-CMG+. Finally, we analyze and validate the simulation platform results, and provide some prospects for future ISAC testing research combined with channel simulators.
@@
ISAC standardization, channel simulator, GBSM, concatenate, ISAC testing.
@@ 1 Introduction
The next-generation mobile communication technology (6G) is expected to advance traditional communication capabilities such as high capacity, low latency, and high reliability while also achieving native intelligence, integrated sensing and communication (ISAC), and ubiquitous connectivity. The boundary between the physical and digital worlds will become increasingly blurred, providing users with a ubiquitous and immersive experience[1, 2]. ISAC technology is anticipated to drive the realization of various use cases, especially in target detection, localization, tracking, imaging, and estimation of target type, speed, and angle. These capabilities will serve applications in autonomous driving, industrial automation, digital twin reconstruction of physical environments, gesture detection, and activity recognition[3, 4, 5]. The sensing information obtained by ISAC can enable the construction of smart homes, smart factories, smart cities, and intelligent transportation systems. Additionally, ISAC technology introduces new performance dimensions compared to traditional communication technologies, such as detection probability, speed, angle, delay, sensing accuracy, and resolution, as well as sensing capacity density[6].
@@
The transmission medium at both the transmitter and receiver determines the performance limits of wireless communication systems. The wireless channel is a critical component in the research and evaluation of wireless communication technologies[7].The evaluation scenarios for ISAC present new key requirements for channel models in terms of sensing modes, target modeling, and modeling methods. Firstly, in terms of sensing modes, traditional channel models are bi-static channels for communication. Sensing scenarios require consideration of mono-static channel models for self-sensing, thus ISAC channel models need to consider both mono-static and bi-static sensing channel models. Secondly, in target modeling for sensing, ISAC applications add the evaluation of target detection, localization, identification, and tracking, such as intrusion detection, UAV and V2X tracking, and recognition of different types and sizes of targets[8]. This necessitates leveraging the scattering properties, polarization characteristics, and Doppler information of electromagnetic waves interacting with targets. Therefore, channel models need to incorporate the electromagnetic scattering characteristics of different targets. Lastly, in modeling methods, it is necessary to consider the cascaded characteristics of channels with targets and the coupling with traditional communication channels[9, 10, 11, 12].
@@
Currently, the standardization of ISAC channel models by 3GPP has reached an intermediate stage, defining ISAC scenarios, sensing targets, sensing modes, and channel model frameworks. The scenarios include traditional UMi, UMa, RMa, InH scenarios, as well as V2X, UAV, and industrial IoT scenarios. According to the scenarios, sensing targets are categorized into five types: humans, vehicles, UAVs, AGVs, and obstacles. Sensing modes are divided into Bi-Static and Mono-Static, and further subdivided into six sensing use cases based on whether the sensing device is a terminal or a base station: BS-UT-Bi, BS-BS-Bi, UT-UT-Bi, UT-BS-Mono, UT-UT-Mono, and BS-BS-Mono[13]. In the channel model framework, the ISAC channel is defined as the sensing channel plus the background channel. The sensing channel includes all parts of the channel interacting with targets, while the background channel includes all other parts of the channel that do not interact with targets. Based on the current progress of 3GPP standardization work and the GBSM simulation framework, this paper proposes an ISAC channel simulation framework aimed at future standardization and integrates it into the BUPTCMCC-6G-CMG+ channel simulation platform. The key characteristics of ISAC channels, such as cascaded characteristics, RCS polarization, scattering properties, and the coupling of target and background channels, are simulated and analyzed to provide a reference model system for ISAC channel simulation.
@@ 2 Simulation Architecture
This paper extends the 3GPP Geometry-Based Stochastic Model (GBSM) channel model by integrating new characteristics of sensing targets, thereby enhancing its scalability while maintaining the continuity of 5G channel models. This extension addresses the simulation requirements for evaluating the sensing performance of ISAC channels. We propose a novel ISAC channel model simulation architecture based on a cascaded channel approach. In this model, the ISAC channel is divided into two components: the target channel and the background channel. The geometric schematic of the model is shown in Figure 1.
The target channel is subdivided into Tx-Target and Target-Rx sub-channels.In the bi-static sensing mode, the target channel is formed by cascading the Tx-Target and Target-Rx sub-channels, with each sub-channel’s parameters generated using the 3GPP statistical model.In the mono-static sensing mode, the Tx-Target link is generated using the statistical model, while the Target-Rx part is derived utilizing channel reciprocity.
As for background channel in the bi-static sensing mode, the background channel is directly generated using the 3GPP statistical channel model.In the mono-static sensing mode, the background channel can either be evolved from the 3GPP method (where related work is limited) or extracted through additional measurements to obtain the channel model parameters.
It is crucial to consider the multipath power distribution characteristics between the target channel and the background channel. The power of both the target and background channels must be normalized to ensure that the overall channel conforms to the 3GPP loss parameters.
@@
Our model framework is an extension of the 3GPP TR38.901 GBSM model[9]. Fig. 2 illustrates the overall simulation process. Users need to perform general configuration in step 1 (i.e., BS, UE, and target positions, antenna configuration, scene, target type). Steps 2-5 pertain to large-scale parameter generation, steps 6-12 focus on small-scale parameter generation, and steps 13-14 generate the channel coefficients.
In particular, steps 4, 9, and 12, which correspond to large-scale parameter cascading, RCS generation, and small-scale parameter cascading, are the extended steps to adapt to ISAC channel characteristics. It is important to note that the steps not highlighted in gray are consistent with the GBSM model generation process. However, each step needs to consider the channel parameter generation for the Tx-Target, Target-Rx, and background channels.
In step 4, the large-scale parameter cascading involves power normalization for the target and background channels. In step 12, the small-scale parameter cascading includes cluster reduction to decrease the simulation complexity.
@@ 2.1 Concatenate channel modeling
We propose a parameter cascading method for ISAC channel simulation based on the scenario layouts, path loss formulas, LOS probability, and large-scale parameters provided in the 3GPP channel model standards TR38.901, TR36.777, and TR37.885. Specifically, TR38.901 is utilized for human and AGV sensing target scenarios, TR36.777 for UAV target scenarios, and TR37.885 for V2X target scenarios. The proposed cascading simulation framework can be divided into large-scale parameter cascading and small-scale parameter cascading, which will be detailed in the following two subsections[14].
@@ 2.1.1 Large scale parameters concatenate
As illustrated in Fig. X, the simulation begins by setting up the simulation scenario, target type, target position, target movement speed and direction, number and location of base stations and terminals, antenna patterns and orientation angles of base stations and terminals, system center frequency, and terminal movement speed and direction. For the configured scenario, the LOS probability and path loss for the Tx-Target and Target-Rx links are calculated separately. Based on the target type, the corresponding RCS mean is generated, and combined with the path loss from both links, the cascaded path loss of the target channel can be written as
@@
Here, $d_{1}$ and $d_{2}$represents the distance from the transmitter to the target and the distance from the target to the receiver respectively; $c$ is the center frequency of the simulation,commonly taken as $3\times 10^{8}$m/s; $\overline{\sigma}_{RCS}$ denotes the mean radar cross-section area of the target.
@@ 2.1.2 Small scale parameters concatenate
The overall channel parameter generation expression used by the BUPTCMCC-6G-CMG+ channel simulator is as follows:
@@
Considering that the cascaded link between the transmitter-target (tx-tar) and target-receiver (tar-rx) involves four possible scenarios as: $LOS+LOS$, $LOS+NLOS$, $NLOS+LOS$, and $NLOS+NLOS$. Equation 2.1.2 presents the general channel expression under the NLOS-NLOS condition. When LOS conditions are involved, the corresponding LOS components in Equation (1), such as power $P_{n}$, angles $\theta$ and $\phi$, and cross-polarization power ratio, should be selectively modified according to the expressions for the LOS case in the GBSM model. In step 4 of Fig. X, large-scale parameters are generated, producing delay, cluster power, angle, XPR, and other small-scale parameters for both the Tx-Target and Target-Rx links. The cascaded small-scale parameters are generated following the process outlined below. Clusters are pruned from the Tx-Target and Target-Rx links based on a predefined power threshold. Let $Q$ and $P$ be the number of clusters remaining in the Tx-Target and Target-Rx links, respectively, after pruning. The total number of clusters in the target channel is $N=QP$. For the n-th cluster of the target channel where $n=P$($q-1$)$+q$.
@@
where, $\tau_{p}$ and $\tau_{q}$ represent the delay of p-th and q-th cluster from Tx-Target and Target-Rx links respectively.
@@
Here, $P_{p}$ and $P_{q}$ represent the power of p-th and q-th cluster respectively.
@@
$\begin{bmatrix}exp(j\Phi_{n}^{\theta\theta})&\sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n}^{\theta\phi})\\ \sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n}^{\phi\theta})&exp(j\Phi_{n}^{\phi\phi})\end{bmatrix}=\begin{bmatrix}exp(j\Phi_{q}^{\theta\theta})&\sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\theta\phi})\\ \sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\phi\theta})&exp(j\Phi_{n}^{\phi\phi})\end{bmatrix}S\\ \begin{bmatrix}exp(j\Phi_{p}^{\theta\theta})&\sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\theta\phi})\\ \sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\phi\theta})&exp(j\Phi_{p}^{\phi\phi})\end{bmatrix}$ (11)
In the 3GPP RAN 1#116 meeting, RAN 1 defined the ISAC channel is composed of a component of target channel and a component of background channel, but how to generate the combined ISAC channel are considered for further study. The relationship between the component of background channel and the environment channel with no sensing target is still an open issue. we propose a new parameter named coupling factor to combine the path loss of target and background channels, and then the path loss of the ISAC channel can be expressed as:
@@
where the $PL_{ISAC}$ and $PL_{target}$ represent the linear path loss values of the ISAC channel and target channel, respectively. The $PL_{backgraound}^{standardization}$ represents the environment channel, i.e., the initial background channel with no sensing target, and it is important to note that $PL_{backgraound}^{standardization}$ is not the same as the defined $PL_{backgraound}$he latter represents the path loss of the channel components after removing the target channel components. Considering the similarity and compatibility with the communication channel, it is suggested to reuse the 3GPP TR 38.901 path loss. The values of $O_{isac}$ needs for further study with more measurement results.
@@ 2.2 RCS modeling
Radar utilizes the scattered power of targets for sensing, and the modeling of the Radar Cross Section (RCS) of targets directly impacts the performance evaluation of sensing. First, the concept of the ”three-dimensional resolution cell” needs to be introduced. A resolution cell is defined as the volume instantaneously illuminated and scattered, comprising two dimensions: angle and distance, and is defined as follows:
@@
where $\frac{c\tau}{2}$ represents the distance resolution cell, $c$ is the speed of light, and $\tau$ is the pulse width; $\Omega$ donates be the antenna beam solid angle, and $R$ is the distance from the radar receiver to the target. If the target’s size is contained within the volume $V$, the target is considered a point target. For the five types of sensing targets defined by 3GPP UAV, Human, Vehicle at a long distance, AGV, and small road obstacles—modeling as point targets is appropriate. For targets larger than the resolution cell and with irregular shapes, such as vehicles and AGVs at close range, large obstacles, and humans in gesture recognition scenarios, multi-point RCS modeling may be necessary.
Table 1 provides sample RCS values (average values) for common targets in the microwave frequency band.

### v2

@@ Abstract
Integrated Sensing and Communication (ISAC) represents a pivotal technology in 6G, with ongoing research and standardization efforts advancing robustly. This paper presents a design and implementation methodology for ISAC channel simulation, utilizing a Geometry-Based Stochastic Model (GBSM) framework. First, we introduce the progress of 3GPP ISAC channel standardization and the key topics of discussion. Second, in response to the existing gap in standardized ISAC channel simulation frameworks, we propose a cascaded ISAC channel simulation framework founded on the Geometry-Based Stochastic Model (GBSM), utilizing our team’s relevant measurements, analyses, and proposed results. Building upon this framework, we have developed and designed the ISAC channel simulator, BUPTCMCC-6G-CMG+. Finally, we analyze and validate the results of the simulation platform, and present prospects for future ISAC testing research in conjunction with channel simulators.
@@
Channel simulator, GBSM, concatenate, ISAC testing, ISAC standardization.
@@ 1 Introduction
Next-generation mobile communication technology (6G) is anticipated to enhance traditional communication capabilities, including high capacity, low latency, and high reliability, while also achieving native intelligence, ISAC, and ubiquitous connectivity. The boundary between the physical and digital realms is expected to become increasingly indistinct, thereby offering users a ubiquitous and immersive experience [1, 2]. ISAC technology is expected to facilitate the implementation of diverse use cases, particularly in target detection, localization, tracking, imaging, and the estimation of target type, speed, and angle. These capabilities will enable applications in autonomous driving, industrial automation, digital twin reconstruction of physical environments, gesture detection, and activity recognition. The sensing information derived from ISAC can facilitate the development of smart homes, factories, smart cities, and intelligent transportation systems. Additionally, ISAC technology introduces new performance dimensions relative to traditional communication technologies, including detection probability, speed, angle, delay, sensing accuracy, resolution, and sensing capacity density[6].
@@
The transmission medium at both the transmitter and receiver establishes the performance limits of wireless communication systems. Furthermore, the wireless channel serves as a critical component in the research and evaluation of wireless communication technologies[7]. The evaluation scenarios for ISAC introduce new key requirements for channel models, particularly regarding sensing modes, target modeling, and modeling methodologies. Firstly, regarding sensing modes, traditional channel models typically represent bi-static channels used for communication purposes. Sensing scenarios necessitate the use of mono-static channel models for self-sensing; therefore, ISAC channel models must incorporate both mono-static and bi-static sensing configurations. Secondly, in the context of target modeling for sensing, ISAC applications expand the evaluation criteria to include target detection, localization, identification, and tracking. Examples include intrusion detection, tracking of UAVs and V2X communications, as well as recognition of various types and sizes of targets. This necessitates the utilization of scattering properties, polarization characteristics, and Doppler information of electromagnetic waves as they interact with targets. Consequently, channel models must integrate the electromagnetic scattering characteristics of various targets. Finally, in modeling methods, it is imperative to account for the cascaded characteristics of channels interacting with targets, as well as their coupling with traditional communication channels.[9, 10, 11, 12].
@@
At present, the standardization of ISAC channel models by 3GPP has reached an intermediate stage, wherein ISAC scenarios, sensing targets, sensing modes, and channel model frameworks have been defined. The defined scenarios encompass traditional Urban Micro (UMi), Urban Macro (UMa), Rural Macro (RMa), and Indoor Hotspot (InH) scenarios, in addition to Vehicle-to-Everything (V2X), Unmanned Aerial Vehicle (UAV), and Industrial Internet of Things (IoT) scenarios. Based on the defined scenarios, sensing targets are classified into five categories: humans, vehicles, unmanned aerial vehicles (UAVs), automated guided vehicles (AGVs), and obstacles. Sensing modes are categorized into bi-static and mono-static configurations, which are further subdivided into six specific use cases based on whether the sensing device is a user terminal (UT) or a base station (BS): BS-UT-Bi, BS-BS-Bi, UT-UT-Bi, UT-BS-Mono, UT-UT-Mono, and BS-BS-Mono.[13]. In the framework of channel modeling, the ISAC channel is defined as the combination of the sensing channel and the background channel. The sensing channel encompasses all components of the channel that interact with targets, whereas the background channel comprises all other components that do not engage with targets.Building upon the current advancements in 3GPP standardization efforts and the GBSM simulation framework, this paper proposes an ISAC channel simulation framework designed for future standardization, which is integrated into the BUPTCMCC-6G-CMG+ channel simulation platform. The key characteristics of ISAC channels—including cascaded features, RCS polarization, scattering properties, and the coupling between target and background channels—are simulated and analyzed to establish a reference model system for ISAC channel simulation.
@@ 2 Simulation Architecture
This paper expands the 3GPP Geometry-Based Stochastic Model (GBSM) channel model by incorporating new characteristics of sensing targets, thereby improving its scalability while preserving the continuity of 5G channel models. This expansion fulfills the simulation requirements for assessing the sensing performance of ISAC channels. We propose an innovative ISAC channel model simulation architecture grounded in a cascaded channel framework. In this model, the ISAC channel is classified into two components: the target channel and the background channel. The geometric schematic of the model is shown in 1.
The target channel is further classified into Tx-Target and Target-Rx sub-channels. In the bi-static sensing mode, the target channel is constructed by cascading the Tx-Target and Target-Rx sub-channels, with parameters for each sub-channel generated according to the 3GPP statistical model. In the mono-static sensing mode, the Tx-Target link is generated using the statistical model, whereas the Target-Rx component is derived by utilizing channel reciprocity.
In the bi-static sensing mode, the background channel is directly generated using the 3GPP statistical channel model. In the mono-static sensing mode, the background channel can be derived from the 3GPP method, which has limited related work, or extracted through additional measurements to obtain the necessary channel model parameters.
Considering the multipath power distribution characteristics between the target channel and the background channel is essential. Both the target and background channel powers must be normalized to ensure that the overall channel adheres to the 3GPP loss parameters.
@@
Our model framework extends the 3GPP TR38.901 GBSM model[9]. Fig. 2 depicts the overall simulation process. In Step 1, users are required to perform general configurations, including the positions of the base station (BS), user equipment (UE), and targets, as well as antenna configuration, scene settings, and target types. Steps 2 to 5 are dedicated to large-scale parameter generation, while Steps 6 to 12 concentrate on small-scale parameter generation. Finally, Steps 13 and 14 are responsible for generating the channel coefficients.
Specifically, Steps 4, 9, and 12, which correspond to large-scale parameter cascading, radar cross-section (RCS) generation, and small-scale parameter cascading, are extended to accommodate the characteristics of ISAC channels. It is essential to emphasize that the steps not highlighted in gray align with the GBSM model generation process. However, each step must take into account the generation of channel parameters for the Tx-Target, Target-Rx, and background channels.
In step 4, the cascading of large-scale parameters entails the normalization of power for both the target and background channels. In step 12, the cascading of small-scale parameters incorporates cluster reduction to mitigate simulation complexity.
@@ 2.1 Concatenate channel modeling
We propose a parameter cascading method for ISAC channel simulation, grounded in the scenario layouts, path loss formulas, line-of-sight (LOS) probabilities, and large-scale parameters delineated in the 3GPP channel model standards TR38.901, TR36.777, and TR37.885. Specifically, TR38.901 is employed for scenarios involving human and automated guided vehicle (AGV) sensing targets, TR36.777 is designated for unmanned aerial vehicle (UAV) target scenarios, and TR37.885 is applied to vehicle-to-everything (V2X) target scenarios. The proposed cascading simulation framework comprises two main components: large-scale parameter cascading and small-scale parameter cascading, which will be elaborated upon in the subsequent two subsections.[14].
@@ 2.1.1 Large scale parameters concatenate
As illustrated in Fig.2, the simulation commences with the establishment of various parameters, including the simulation scenario, target type, target position, movement speed and direction of the target, the number and location of base stations and terminals, antenna patterns, orientation angles of both base stations and terminals, system center frequency, and movement speed and direction of the terminals. For the configured scenario, the Line-of-Sight (LOS) probability and path loss for the Tx-Target and Target-Rx links are computed individually. Based on the target type, the corresponding Radar Cross Section (RCS) mean is generated. This value, when combined with the path loss from both links, allows the cascaded path loss of the target channel to be expressed as
@@
In this context, $d_{1}$ and $d_{2}$ represent the distances from the transmitter to the target and from the target to the receiver, respectively. Here, $c$ denotes the center frequency of the simulation, typically considered as $3\times 10^{8}$ m/s, while $\overline{\sigma}_{RCS}$ indicates the mean radar cross-section area of the target.
@@ 2.1.2 Small scale parameters concatenate
The overall expression for channel parameter generation utilized by the BUPTCMCC-6G-CMG+ channel simulator is as follows:
@@
Considering that the cascaded link between the transmitter-target (tx-tar) and target-receiver (tar-rx) involves four possible scenarios: $LOS+LOS$, $LOS+NLOS$, $NLOS+LOS$, and $NLOS+NLOS$. Equation 2.1.2 presents the general channel expression under the NLOS-NLOS condition. When LOS conditions are involved, the corresponding LOS components in Equation (1), such as power $P_{n}$, angles $\theta$ and $\phi$, and cross-polarization power ratio, should be selectively modified according to the expressions for the LOS case in the GBSM model. In Step 4 of Fig. 2, Large-scale parameters are generated, resulting in the production of delay, cluster power, angles, XPR, and other small-scale parameters for both the Tx-Target and Target-Rx links. Subsequently, the cascaded small-scale parameters are generated in accordance with the outlined process. Clusters are pruned from the Tx-Target and Target-Rx links based on a predefined power threshold. Let $Q$ and $P$ be the number of clusters remaining in the Tx-Target and Target-Rx links, respectively, after pruning. The total number of clusters in the target channel is $N=QP$. For the n-th cluster of the target channel where $n=P$($q-1$)$+q$.
@@
where $\tau_{p}$ and $\tau_{q}$ represent the delay of the p-th and q-th cluster from Tx-Target and Target-Rx links respectively.
@@
Here, $P_{p}$ and $P_{q}$ represent the power of the p-th and q-th cluster respectively.
@@
Cross polarization power ratio (XPR):
$XPR=\begin{bmatrix}exp(j\Phi_{q}^{\theta\theta})&\sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\theta\phi})\\ \sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\phi\theta})&exp(j\Phi_{n}^{\phi\phi})\end{bmatrix}S\\ \begin{bmatrix}exp(j\Phi_{p}^{\theta\theta})&\sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\theta\phi})\\ \sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\phi\theta})&exp(j\Phi_{p}^{\phi\phi})\end{bmatrix}$ (11)
During the 3GPP RAN 1 #116 meeting, RAN 1 defined the ISAC channel as comprising a component from the target channel and a component from the background channel. However, the methodology for generating the combined ISAC channel is subject to further study. The relationship between the background channel component and the environment channel in the absence of sensing targets remains an unresolved issue. To address this, we introduce a new parameter, termed the coupling factor, to integrate the path loss of both the target and background channels. Consequently, the path loss of the ISAC channel can be expressed as follows:
@@
where the $PL_{ISAC}$ and $PL_{target}$ represent the linear path loss values of the ISAC channel and target channel, respectively. The $PL_{background}^{standardization}$ represents the environment channel, i.e., the initial background channel with no sensing target, and it is important to note that $PL_{background}^{standardization}$ is not the same as the defined $PL_{backgraound}$ The latter refers to the path loss of the channel components after the removal of the target channel components. Given the similarities and compatibility with the communication channel, it is advisable to reuse the path loss model from 3GPP TR 38.901. The values of $O_{isac}$ need for further study with more measurement results.
@@ 2.2 RCS modeling
Radar systems leverage the scattered power of targets for sensing, and the accurate modeling of the Radar Cross Section (RCS) of these targets directly influences the performance evaluation of sensing capabilities. First, it is essential to introduce the concept of the ”three-dimensional resolution cell.” A resolution cell is defined as the volume that is instantaneously illuminated and scattered, characterized by two dimensions: angle and distance. It can be formally expressed as follows:
@@
where $\frac{c\tau}{2}$ represents the distance resolution cell, $c$ is the speed of light, and $\tau$ is the pulse width; $\Omega$ donates be the antenna beam solid angle, and $R$ is the distance from the radar receiver to the target. If the target’s size is contained within the volume $V$, the target is considered a point target. For the five types of sensing targets defined by 3GPP—UAVs, humans, long-distance vehicles, AGVs, and small road obstacles—modeling as point targets is appropriate. However, for targets that exceed the resolution cell and possess irregular shapes, such as close-range vehicles and AGVs, large obstacles, and humans in gesture recognition scenarios, multi-point RCS modeling may be required.